OpenStudy (anonymous):

For a triangle with sides a b c 3(bc+ca+cb) <= (a+b+c)^2 <=4(bc +ca+ab)

OpenStudy (anonymous):

and the Question would be?

OpenStudy (anonymous):

Show that it is so.

OpenStudy (sriram):

[a+b+c]^2=a^2+b^2+c^2+2(ab+bc+ca) hence to prove (bc+ca+ab) <= a^2+b^2+c^2 <=2(bc +ca+ab) it is quite clear that (bc+ca+ab) <=2(bc +ca+ab) now we need to prove that a^2+b^2+c^2 lies btwn the two

OpenStudy (anonymous):

Good start...

OpenStudy (sriram):

AM>=GM 4ab<=(a+b)^2 a^2+b^2>=2ab similarly b^2+c^2>=2bc c^2+a^2>=2ca adding the three we get 2[a^2+b^2+c^2]>=2(ab+bc+ca) a^2+b^2+c^2] >= ab+bc+ca